# LLM2SMT: Building an SMT Solver with Zero Human-Written Code

### 06 Mar 2026 v1

@@ Abstract
Whether LLMs can reason or write software is widely debated, but whether they can write software that itself reasons is largely unexplored. We present a case study in which an LLM coding agent builds a complete DPLL(T)-style SMT solver for QF_UF with zero human-written code. The solver implements the Nieuwenhuis-Oliveras congruence closure algorithm, includes preprocessing, and emits Lean proofs for unsatisfiable instances. We describe the development process and key challenges, and show that the resulting solver is competitive on SMT-LIB benchmarks.
@@ 2 Background
Lean [21] is a popular modern interactive theorem prover (ITP) and programming language for formalizing mathematics and verifying software. It is based on dependent type theory, providing a unified framework for programs, proofs, and mathematical objects. Its main library, mathlib, contains thousands of formalized theorems spanning broad areas of mathematics.
@@ 3.3 Certification
The correctness of the proof mechanism was probably the most challenging task of this case study. Repeated fuzzing and bug fixing were required for the agent to correctly implement all steps of the proof. Another obstacle is that the agent does not grasp the complexity at the Lean proof level. We had trouble communicating to the agent that only easy steps should be presented to the grind tactic. Table˜2, in the upcoming section, presents the experimental results on the whole benchmark set.
@@ 6 Conclusion and Future Work
The central question we posed was: Can LLMs develop automated reasoning tools? The answer is a qualified yes. With appropriate scaffolding—systematic fuzzing, delta-debugging, explicit resource limits, and careful specification of input formats—a coding agent can produce a competitive SMT solver from scratch, solving nearly as many benchmarks as mature solvers such as z3 and cvc5—admittedly on a single theory. The agent is capable of implementing techniques described in scientific articles and debugging its own code when given concrete failing examples.
Two important caveats emerged. First, correctness cannot be assumed: the agent requires guidance, concrete examples, and systematic testing throughout; subtle errors (such as mishandling Booleans as both propositions and terms) can persist unnoticed without a fuzzer. Second, proof generation proved to be the hardest challenge: the agent must understand not just the solver logic but also the expectations of the proof checker (Lean in our case), which required significant human guidance.
In the future we aim to add more functionality to the solver but also prefer more careful analysis of the written code to suggest improvements to the agent.

### 10 Mar 2026 v2

@@ Abstract
Whether LLMs can reason or write software is widely debated, but whether they can write software that itself reasons is largely unexplored. We present a case study in which an LLM coding agent builds a complete DPLL(T)-style SMT solver for QF_UF with zero human-written code. The solver implements the Nieuwenhuis-Oliveras congruence closure algorithm, includes preprocessing, and emits Lean proofs for unsatisfiable instances. We describe the development process and key challenges, and show that the resulting solver is competitive on SMT-LIB benchmarks. While full proof emission for unsatisfiable instances appears to be beyond the capabilities of the LLM, the solver achieves competitive performance and passes correctness tests.
@@ 2 Background
Lean [21] is a popular modern interactive theorem prover (ITP) and programming language for formalizing mathematics and verifying software. It is based on dependent type theory, providing a unified framework for programs, proofs, and mathematical objects. Its main library, mathlib, contains a bevy of formalized theorems spanning broad areas of mathematics.
@@ 3.3 Certification
The correctness of the proof mechanism was probably the most challenging task of this case study. Repeated fuzzing and bug fixing were required for the agent to correctly implement all steps of the proof. Another obstacle is that the agent does not grasp the complexity at the Lean proof level. We had trouble communicating to the agent that only easy steps should be presented to the grind tactic. At the same time, the agent seems not to be able realize which steps need to be encoded into the proof. Table˜2, in the upcoming section, presents the experimental results on the whole benchmark set.
@@ 6 Conclusion and Future Work
The central question we posed was: Can LLMs develop automated reasoning tools? The answer is a qualified yes. With appropriate scaffolding—systematic fuzzing, delta-debugging, explicit resource limits, and careful specification of input formats—a coding agent can produce a competitive SMT solver from scratch, solving nearly as many benchmarks as mature solvers such as z3 and cvc5—admittedly on a single theory. The agent is capable of implementing techniques described in the research literature and debugging its own code when provided with concrete failing examples.
Directing the agent toward established techniques such as delta-debugging and fuzzing proved beneficial. Proof generation was the most challenging aspect: the agent must understand not only the solver logic but also the expectations of the proof checker (Lean in our case), which required significant human guidance and still fails on a number of problems. The agent does grasp the structure of proof construction—in several cases it successfully corrected its own invalid proofs—yet it struggles to faithfully encode all computational information produced during solving into the final proof certificate.
In future work we plan to extend the solver with additional theories and to incorporate more systematic human code review into the agent’s feedback loop, with the goal of identifying and guiding improvements to the generated implementation (human-in-the-loop).